What is correct ans plz tell , 
√36 = +-6 , +6 ,- 6 ????

Grade:10

1 Answers

Vijay
2 years ago
By mathematical defination it is given that – 
\sqrt{ x^{2} } = \left | X \right |  
so , 
 
 \sqrt{ 6^{2} } = \left | 6 \right |   
 
But when we get something like that x^{2} = 36 \Rightarrow x = \pm 6 is the answer 
 
I hope You understand 
Thanks